Find the equation of the line containing the point (-5, -8) and slope m = -3
Found 2 solutions by math_helper, ikleyn:
Answer by math_helper(2461)


y - (-8) = -3 (x - (-5))
y + 8 = -3(x + 5) point slope form
y + 8 = -3x - 15

y = -3x - 23 slope intercept form
3x + y = -23 standard form

Answer by ikleyn(52781)

.
Learn the subject from the lessons
    - Find the slope of a straight line in a coordinate plane passing through two given points
    - Equation for a straight line having a given slope and passing through a given point (*)
    - Solving problems related to the slope of a straight line
    - Equation for a straight line in a coordinate plane passing through two given points
    - Equation for a straight line parallel to a given line and passing through a given point
    - Equation for a straight line perpendicular to a given line and passing through a given point
in this site.

The most relevant lesson is marked by (*) in the list - so start from it.

But if you want to learn the subject in whole, then look into all these lessons.

Consider them as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.